What is the equivalent present amount of an eight-year series of decreasing amounts if the interest rate is 12% compounded annually, the first year amount is $35,000, and the rate of decrease is $1050 per year?

2-29 na Base Amout AT 20,000 -ve 800 ACPIAI A-G CA16i,n)) (eLAi,n)| (A1,0)(P1 800 20,000 3-0 045 S-33 49 93,875-03

I was giving this example but I dont know how to get $93875.03,

I did 20000 - 800(3.0045)(5.3349)= 7177.03, what am I doing wrong?
Answer #1

YOU WRITING IN A WRONG WAY. BRACKETS ARE PLACED IN A WRONG WAY

IT IS : (20000 - 800*(3.0045)) * (5.3349) = (20000 - 2403.60) * (5.3349) = (17596.40) * (5.3349) = 93785.03

SAME WAY, WE CAN DO THE GIVEN SUM

(A/G, 12%, 8) = 2.9131, (P/A, 12%, 8) = 4.9676

IT WILL BE : (35000 - 1050*(2.9131))*(4.9676) = (31941.245)*(4.9676) = 158671.33

ANSWER : 158671.33 (Thumbs up please) [ANY DOUBTS FEEL FREE TO ASK]